Techuxia Corporation worked on four jobs during October: Job A256, Job A257, Job A258, and Job A260. At the end of October, the job cost sheets for these jobs contained the following data:   

Job A256 Job A257 Job A258 Job A260
Beginning balance $ 1,000 $ 660 $ 0 $ 0
Charged to the jobs during October:
Direct materials $ 2,800 $ 4,020 $ 1,580 $ 3,790
Direct labor $ 1,170 $ 910 $ 620 $ 320
Manufacturing overhead applied $ 3,976 $ 1,760 $ 1,934 $ 420
Units completed 213 0 106 0
Units in process at the end of October 0 304 0 257
Units sold during October 157 0 58 0

Jobs A256 and A258 were completed during October. The other two jobs had not yet been completed at the end of October. There was no finished goods inventory on October 1.

       

In October, overhead was overapplied by $1,350. The company adjusts its cost of goods sold every month for the amount of the underapplied or overapplied overhead.

Required:     

1. Using the direct method, what is the cost of goods sold for October?

2. What is the total value of the finished goods inventory at the end of October?

      

3. What is the total value of the work in process inventory at the end of October?
Answer #1

Cost per unit JobA256 = ($1,000 + $2,800 + $1,170 + $3,976) / 213 = $42

Cost per unit JobA258 = ($1,580 + $620 + $1,934) / 106 = $39

1.

Cost of goods sold = (157 X $42) + (58 X $39) - $1,350 = $7,506

Finished goods JobA256= 213 - 157 = 56

Finished goods JobA258 = 106 - 58 = 48

2.

Finished goods = (56 X $42) + (48 X $39) = $4,224

3.

Total value of work in process = Cost of JobA257 & JobA260

= $660 + $4,020 + $910 + $1,760 + $3,790 + $320 + $420

= $11,880